OpenStudy (anonymous):

Find lim x->2 of (x^3 - 8)/(x^4 -16)

OpenStudy (anonymous):

You did this one before I believe where the top and bottom both came out to zero....Do you remember what you ended up doing to solve it?

OpenStudy (anonymous):

nope....no idea....

OpenStudy (anonymous):

Something called....L'Hopital rule perhaps?

OpenStudy (anonymous):

i believe you can factor the numerator and denominator

OpenStudy (anonymous):

find the limit algebraically, not using L'hopitals rule

OpenStudy (anonymous):

Yep, you can do that too

OpenStudy (anonymous):

maybe get rid of the (x - 2) term

OpenStudy (anonymous):

factorize numerator and denominator and cancel the terms. you leave be left be some polynomials. substitute x=2 and you will get the answer.

OpenStudy (anonymous):

the numberator is a difference of 2 cubes, numerator is a difference of 2 squares (twice)

OpenStudy (anonymous):

*you will be left with some polynomials.

OpenStudy (anonymous):

numberator* lol numerator

OpenStudy (anonymous):

or if you want to do it algebraically, the top factors to (x-2)(x^2+2x+2)

OpenStudy (anonymous):

the bottom factors to (x^2+4)(x-2)(x+2)

OpenStudy (anonymous):

cancelling you have (x^2+2x+2)/(x^2+4)(x+2)

OpenStudy (anonymous):

the numerator does not factor to (x-2)(x^2+2x+2)

OpenStudy (anonymous):

multiplying those together gives you x^3 - 4x -4

OpenStudy (anonymous):

now since, the zero in the denominator has been cancelled, you can just plug in the values and get whatever

OpenStudy (anonymous):

except that you dont because you didnt factor correctly

OpenStudy (anonymous):

it's (x^2+2x+4)/(x^2+4)(x+2)

OpenStudy (anonymous):

oops, the numerator factors to (x-2)(x^2+2x+4)

OpenStudy (anonymous):

that works...

OpenStudy (anonymous):

you still have a zero-free denominator

OpenStudy (anonymous):

plugging in, you get 12/32, reduced to 3/8

OpenStudy (anonymous):

Yeah, I forgot to multiply

OpenStudy (anonymous):

Yes, correct, it's 3/8

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

tell me this....how do you figure out that x^4 - 16 is equal to (x^2 + 4)(x-2)(x+2) ? What are the steps involved...?

OpenStudy (anonymous):

difference of squares

OpenStudy (anonymous):

(x^2+4)(x^2-4)

OpenStudy (anonymous):

(x^2+4)(x^2-4) = (x^2+4)(x+2)(x-2)

OpenStudy (anonymous):

and then x^2-4 factors to (x+2)(x-2)

OpenStudy (anonymous):

the formula for the difference of squares is: \[a^2-b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

alright.....been awhile since i had to do anything like that.....

OpenStudy (anonymous):

how about the numerator? is that difference of cubes?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

alright, i looked those up....just forgot those equations existed...

OpenStudy (anonymous):

Formula for the difference of cubes is: \[a^3-b^3 = (a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

\[a ^{3} - b ^{3} = (a-b)(a ^{2}+ab-b ^{2})\]